OpenStudy (twizttiez):

Complete the equation for the linear function whose graph contains the points (2, 5) and (3, 0). y − = (x − 2)

OpenStudy (misty1212):

HI!!

OpenStudy (twizttiez):

Hello misty xD

OpenStudy (misty1212):

what a weird question ! the points slope formula is \[y-y_1=m(x-x_1)\]

OpenStudy (misty1212):

you have it all filled in except the \(x_1\) part from the ordered pair \((2,5)\)

OpenStudy (twizttiez):

Ok now I'm awful at math what would my \[x1 \] be?

OpenStudy (misty1212):

lol you tell me dear in the ordered pair \((2,5)\) which is the x and which is the y? hint, they are in alphabetical order

OpenStudy (twizttiez):

Lol ok well i think my x would be 2 and my y would be 5 correct?

OpenStudy (misty1212):

oops i see i made a mistake you have it all filled in but the \(y_1\) part which is 5, right

OpenStudy (misty1212):

drop a 5 in the slot and you are done\[\color\magenta\heartsuit\]

OpenStudy (twizttiez):

So would i do y-5=(3-2)?

OpenStudy (twizttiez):

Ohhhhhhhhhh nvm *facepalm* y-5=3(x-2)?

OpenStudy (twizttiez):

:?

OpenStudy (twizttiez):

@misty1212

OpenStudy (twizttiez):

Ya there?